# Binary-Report Peer Prediction for Real-Valued Signal Spaces

### v1

@@ C.1 Proof of Theorem 9
Corollary 6.
Lemma 1.
Suppose agent $j$ plays a consistent strategy. Let $R^{\prime}_{t}$ be the random variable on $\{H,L\}$ resulting agent $j$’s strategy, i.e., $R_{t}^{\prime}\sim\hat{\sigma}(S_{t}^{\prime})$ for each task $t$. Let $p(s_{t})=\Pr[R_{t}^{\prime}=H\mid S_{t}=s_{t}]$, which by assumption is independent of $t$. Then for agent $i$,
where as before
[…]
@@
As $R_{1}^{\prime},R_{2}^{\prime}$ are independent, we have
We thus have
$=p(s_{1})-p(s_{2})~{}.$
We have now established
To finish the proof, consider the cases $s_{1}=s_{2}$, $(s_{1},s_{2})=(H,L)$, and $(s_{1},s_{2})=(L,H)$. From the above, we obtain 0 in the first case. In the second, we have $p(H)-p(L)$ if $(r_{1},r_{2})=(s_{1},s_{2})$ and its negation when $(r_{1},r_{2})=(s_{2},s_{1})$. In the third, we again have $p(H)-p(L)$ if $(r_{1},r_{2})=(s_{1},s_{2})$ and its negation when $(r_{1},r_{2})=(s_{2},s_{1})$.
Proof of Theorem 9.
$=\mathop{\mathbb{E}}[\det M_{12}\mid S_{1},S_{2}]\mathop{\mathbb{E}}[\det M_{34}\mid S_{3},S_{4}]~{}.$
Applying Lemma 1 to both $M_{12}$ and $M_{34}$, observe that if $s_{1}=s_{2}$ or $s_{3}=s_{4}$ we have $\mathop{\mathbb{E}}[M_{\textrm{DMI}}(r_{1..T},R^{\prime}_{1..T})\mid S_{1..T}]=0$ regardless of $r$. In particular, $\sigma_{\text{true}}$ maximizes the expected payoff.
Appendix D Omitted Proofs for RBTS
D.1 Gaussian Model
We first prove the form of the functions $\bar{P}(\tau;x)$ and $Q(x)$ for the Gaussian model.
@@ D.1 Gaussian Model
Now we have the tools to prove the characterization of equilibria in the Gaussian model.
Proof of Corollary 5.
[…]
@@
Let $\tau$ be the current threshold strategy that all other agents are following. By equations (18) and (19), an agent who receives signals $x$ will best respond with $H$ if $P(\tau;x)\leq\bar{P}(\tau;x)$, and $L$ otherwise. In the Gaussian model, note that
$P(0;x)$ $=\Phi\left(\frac{-\rho x}{b\sqrt{1+\rho}}\right),$
$\bar{P}(0;x)$ $=\Phi\left(\frac{-\rho^{2}}{b\sqrt{(1+\rho^{2})(1+\rho)}}\right).$
Thus $P(0;x)=\bar{P}(0;x)$ at a unique point which can be verified with simple algebra to be $x=\hat{\tau}$. One can check for the specified value of $\hat{\tau}$, $P(\tau;x)\geq\bar{P}(\tau;x)$ for $x\leq\hat{\tau}$ and $P(\tau;x)\leq\bar{P}(\tau;x)$ for $x\geq\hat{\tau}$. The result follows. ∎
Now we formally compare convergence of the DG and RBTS mechanisms under the Gaussian model.
@@
$m_{\textrm{DG}}(\rho)=\frac{a-b\sqrt{p(1+p)}}{a\rho}.$
@@
Now, one can check that $\frac{m_{\textrm{DG}}(\rho)}{m_{\textrm{RBTS}}(\rho)}>1$ for all $a,b>0$. It immediately follows that the best response $\hat{\tau}$ converges to 0 strictly faster under the DG mechanism than under the RBTS mechanism. ∎
@@ Appendix E Omitted Experiments
We include bifurcation figures for the four-component Gaussian mixture case within the DG and OA mechanisms, as well as bifurcation figures for RBTS. Again, we vary the precision of each component in the Gaussian mixture and study how the stability and number of equilibria change across mechanisms.
@@
(a) Bifurcation graphs for OA (left) and DG (right) in the setting of a four-component Gaussian mixture with means ($-4,-2,2,4$). Following a pattern, we have seven equilibria in OA and five in DG for large enough precision.
(a) 3 component mixture with means (-2, 0, 2)
(b) 4 component mixture with means (-3, -1, 1, 3)

### v2

@@ C.1 Equilibrium characterization
Let a finite threshold $\tau$ be given and $P(\tau;x)$ be continuous over $x$; also assume $\beta(x)$ is continuous over $x$. If $\tau\in\mathbb{R}$ is threshold equilibrium under the RBTS mechanism then $Q(\tau)=G(\tau)$. Conversely, if $Q(\tau)=G(\tau)$ and $P(\tau;x)-\bar{P}(\tau;x)$ has a single crossing of 0 from positive to negative then $\tau$ is a threshold equilibrium under the RBTS mechanism.
@@
For necessity, first note that since $\beta(x)$ and $P(\tau;x)$ are continuous over $x$, $\bar{P}(\tau;x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}P(\tau;x^{\prime})$ is also continuous over $x$ by an application of the Dominated Convergence Theorem. Let $f(\tau;x)=P(\tau;x)-\bar{P}(\tau;x)$. Since $f(\tau;x)$ is continuous, we must have $\lim_{x\to\tau^{+}}f(\tau;x)=\lim_{x\to\tau^{-}}f(\tau;x)=f(\tau;\tau).$ Further, $\lim_{x\to\tau^{-}}f(\tau;x)\geq 0$ and $\lim_{x\to\tau^{+}}f(\tau;x)\leq 0$ by Conditions 16 and 17, so $f(\tau;\tau)=0=G(\tau)-Q(\tau)$.
While the exact crossing condition has changed from $G(\tau)=F(\tau)$ to $G(\tau)=Q(\tau)$, Theorem 7 paints a similar picture to what we saw for DG. Best responses are now more complex to compute, so we do not treat the general dynamics formally, but we can provide a comparison of their behavior in the Gaussian case.
C.2 Gaussian model
We return to our Normal example, where each agent receives a noisy version of a Gaussian signal. We first prove the form of the functions $\bar{P}(\tau;x)$ and $Q(x)$ in this setting.
@@ C.2 Gaussian model
We can then show that the same equilibria occur in RBTS as the previous mechanisms.
Proposition 11.
In the Gaussian model under RBTS, there are three equilibria at $\tau=0$ and $\pm\infty$.
Proof.
Existence of equilibria at $\pm\infty$ follow as a corollary from 5. Next, note all non-infinite equilibria must satisfy $G(x)=Q(x)$ by Theorem 7. But $G(x)$ and $Q(x)$ both correspond to the standard Normal CDF with different coefficients $c_{G}=\frac{(1-\rho)}{b\sqrt{1+\rho}}$ and $c_{Q}=\frac{(1-\rho^{2})}{b\sqrt{(1+\rho)(1+\rho^{2})}}$. Since $\rho\in(0,1)$,
@@
We can also derive the exact form of the best response $\hat{\tau}=c(\rho)\tau$.
@@
Let $\tau$ be the current threshold strategy that all other agents are following. By equations (16) and (17), an agent who receives signals $x$ will best respond with $H$ if $P(\tau;x)\leq\bar{P}(\tau;x)$, and $L$ otherwise. In the Gaussian model, note by Proposition 11 that
$P(\tau;x)$ $=\Phi\left(\frac{\tau-\rho x}{b\sqrt{1+\rho}}\right),$
Thus $P(\tau;x)=\bar{P}(\tau;x)$ at a unique point which can be verified with simple algebra to be $x=\hat{\tau}$. One can check for the specified value of $\hat{\tau}$, $P(\tau;x)\geq\bar{P}(\tau;x)$ for $x\leq\hat{\tau}$ and $P(\tau;x)\leq\bar{P}(\tau;x)$ for $x\geq\hat{\tau}$. The result follows. ∎
Now, note that for any $\tau$, $\hat{\tau}=c(\rho)\tau$ for $c(\rho)\in(0,1)$. Under our dynamics, $\dot{\tau}=\hat{\tau}-\tau.$ If $\tau>0$, then $\hat{\tau}<\tau$, while if $\tau<0$, $\hat{\tau}>\tau$. It follows that $\tau=0$ is stable, while by topological necessity the uninformative equilibria are unstable.
We observe that the best responses under DG and RBTS are both linear functions of $\tau$ with coefficients $m_{\textrm{DG}}$, $m_{\textrm{RBTS}}$ respectively. Thus we can formally compare convergence of the DG and RBTS mechanisms under the Gaussian model using the best response form derived in Proposition 12. We find that if both mechanisms begin with the same starting threshold $\tau(0)$, the stable equilibrium at 0 is reached faster in DG than RBTS. E.g. in a peer grading setting, graders will more quickly converge to a half-good half-bad quality consensus than RBTS graders, so that the designer may have more time to react and reset expectations in the latter case.
@@
$m_{\textrm{DG}}(\rho)=\frac{a-b\sqrt{\rho(1+\rho)}}{a\rho}.$
@@
Now, one can check that $\frac{m_{\textrm{DG}}(\rho)}{m_{\textrm{RBTS}}(\rho)}<1$ for all $a,b>0$. It immediately follows that the best response $\hat{\tau}$ converges to 0 strictly faster under the DG mechanism than under the RBTS mechanism. ∎
Appendix D Omitted results for DMI
In this section, we characterize equilibria and their dynamics for the DMI mechanism. Unlike the previous mechanisms, where we could compute the payment for a single task in isolation (albeit requiring multiple tasks from a different agent), DMI is inherently a multi-task mechanism. As in the original analysis of DMI, we therefore restrict to consistent strategies, where the same $\sigma$ is applied to each $X_{i}$. We include a discussion of behavior outside consistent strategies in § D.3. Even under this restriction, threshold equilibria are much more complex in DMI given the interactions between four different signals, making it unclear how to reason directly from the definition of a threshold equilibrium. Thus we also restrict all agents to best responding with threshold strategies.
D.1 Equilibrium Characterization
Let
$U_{i}(\sigma,\sigma^{\prime})=\mathop{\mathbb{E}}\left[M_{\textrm{DMI}}(\sigma(X_{1}),\ldots,\sigma(X_{4}),\sigma^{\prime}(X_{1}^{\prime}),\ldots,\sigma^{\prime}(X_{4}^{\prime}))\right]$ (20)
be the (ex-ante) expected utility for playing threshold strategy $\sigma$ given the other agent’s strategy $\sigma^{\prime}$.
Definition 3.
A threshold strategy $\sigma:\mathbb{R}\to\mathcal{R}$ is a equilibrium restricted to threshold strategies under DMI if for all threshold strategies $\hat{\sigma}$$\mathbb{R}\to\mathcal{R}$, $U_{i}(\sigma,\sigma)\geq U_{i}(\hat{\sigma},\sigma)$.
This is a weaker condition than being a threshold equilibrium because it rules out deviations to non-threshold strategies, an issue we will return to. However, by restricting to this space we can provide an equilibrium characterization in the same spirit as we did for OA and DG.
This unbiasedness is implied by Kong [2024, Claim 4.4], which we paraphrase for the binary-report context: for any $Y,Y^{\prime}$, there exists a constant $\alpha>0$ such that $\mathop{\mathbb{E}}\det\sum_{i\in S}\mathbf{1}_{Y}\mathbf{1}_{Y^{\prime}}^{T}=\alpha\det U_{Y,Y^{\prime}}$.
D.2 Equilibrium results
Proposition 14.
$\tau^{*}=\pm\infty$ are both always equilibria restricted to threshold strategies under DMI.
Theorem 8.
Proof.
$0$ $=\frac{d}{d\tau}\alpha^{\prime}h(\tau,\tau^{*})^{2}=2h(\tau,\tau^{*})\alpha^{\prime}f(\tau)\left(\Pr[X^{\prime}>\tau^{*}]-\Pr[X^{\prime}>\tau^{*}\mid X=\tau]\right)$~ (22)
As discussed in the main text, Theorem 8 shows that DMI has the same necessary condition for equilibrium as DG. Thus, beyond the slightly different sufficient condition for equilibrium, its equilibria can be found at the same thresholds: those where $G(\tau)=F(\tau)$. Moreover, with the appropriate sufficient condition, our results about the dynamics (Theorem 5) for DG immediately apply to DMI as well. In particular, the Gaussian case has the same equilibra with the same stability.
D.3 Beyond consistent threshold strategies
The above analysis is restricted to our weaker notion of equilibrium in the space of consistent threshold strategies. It is unclear whether these remain equilibria if agents can use any consistent strategy, or if agents can use any four threshold strategies. Even more broadly, one could consider joint-task strategies $\sigma:\mathbb{R}^{4}\to\{H,L\}^{4}$, which map all four signals simultaneously to all four reports. While not shown in Kong [2020, 2024], in the binary report setting the DMI mechanism turns out to satisfy the even stronger property that truthfulness is an equilibrium among all joint-task strategies.
Specifically, we can show in the binary signal, binary report model that truthfulness remains a best response even when the other agent deviates from truthful, but crucially, still plays a consistent strategy, which here means strategies that (a) map each signal to a distribution over reports, and (b) use the same such map for every task.
Definition 4.
We are able to prove that truthfulness maximizes the expected payoff of $M_{\textrm{DMI}}$ when all other agents play consistent strategies. As a corollary, truthfulness is a Bayes Nash equilibrium even over the larger space of strategies, where one can choose all $T$ reports simultaneously after looking at all $T$ signals. We leave the proof to § D.4.
Theorem 9.
Corollary 5.
We can then show that truthfulness is not an equilibrium for real-valued reports: in many cases the best response flips one or two of the truthful reports to increase the chance of a nonzero payoff in $M_{\textrm{DMI}}$. To get some intuition, first consider the binary signal model and an agent who receives signals $H,L,L,L$. From the form of the mechanism, if they report truthfully, their score will be zero deterministically, since $M_{34}$ will have rank 1. Thus, they might consider deviating, say flipping the last $L$ to $H$. It turns out doing so keeps their expected score at zero, since the other agent is equally likely to align or misalign with their reports. The consistency of the other agent’s strategies is crucial for this last fact; given any slight difference in their strategy for tasks 3 and 4, the first agent would deviate from truthfulness.
Thus, given real-valued signals (10,10,10,1)(10,-10,-10,-1), the optimal report is $(H,L,L,H)$, which is not “truthful” for the prescribed threshold $\tau=0$. Hence DMI is no longer truthful in this joint-task sense, even at the equilibrium threshold $\tau=0$.
D.4 Proof of Theorem 9
Lemma 1.
Suppose agent $j$ plays a consistent strategy. Let $R^{\prime}_{t}$ be the random variable on $\{H,L\}$ resulting from agent $j$’s strategy, i.e., $R_{t}^{\prime}\sim\hat{\sigma}(S_{t}^{\prime})$ for each task $t$. Let $p(s_{t})=\Pr[R_{t}^{\prime}=H\mid S_{t}=s_{t}]$, which by assumption is independent of $t$. Then for agent $i$,
where as before
$M_{12}\;=\;\mathbf{1}{r_{1}}\,\mathbf{1}{R_{1}^{\prime}}^{\top}\;+\;\mathbf{1}{r_{2}}\,\mathbf{1}{R_{2}^{\prime}}^{\top}~.$
Proof.
As $R_{1}^{\prime},R_{2}^{\prime}$ are independent, we have
We thus have
$=p(s_{1})-p(s_{2})~.$
We have now established
To finish the proof, consider the cases $s_{1}=s_{2}$, $(s_{1},s_{2})=(H,L)$, and $(s_{1},s_{2})=(L,H)$. From the above, we obtain 0 in the first case. In the second, we have p(H)p(L)p(H)-p(L) if $(r_{1},r_{2})=(s_{1},s_{2})$ and its negation when $(r_{1},r_{2})=(s_{2},s_{1})$. In the third, we again have p(H)p(L)p(H)-p(L) if $(r_{1},r_{2})=(s_{1},s_{2})$ and its negation when $(r_{1},r_{2})=(s_{2},s_{1})$.
Proof of Theorem 9.
Applying Lemma 1 to both $M_{12}$ and $M_{34}$, observe that if $s_{1}=s_{2}$ or $s_{3}=s_{4}$ we then have $\mathop{\mathbb{E}}[M_{\textrm{DMI}}(r_{1..T},R^{\prime}_{1..T})\mid S_{1..T}]=0$ regardless of $r$. In particular, $\sigma_{\text{true}}$ maximizes the expected payoff.
@@ Appendix E Omitted Experiments
We include bifurcation figures for the four-component Gaussian mixture case within the DG and OA mechanisms (Figure 9(a)), as well as bifurcation figures for RBTS (Figure 9(b)). Again, we vary the precision of each component in the Gaussian mixture and study how the stability and number of equilibria change across mechanisms.
@@
(a) Bifurcation graphs for OA (left) and DG (right) in the setting of a four-component Gaussian mixture with means (4,2,2,4-4,-2,2,4). Following a pattern, we have seven equilibria in OA and five in DG for large enough precision.